### v1

@@ III-B LLM Pre-training
The aim of pre-training language models is to predict the next word within a sentence. After being trained on extensive datasets, LLM models exhibit emergent capabilities in comprehension and reasoning. This subsection will introduce dataset collection, preprocessing, and model training techniques.
@@ III-B1 Dataset collection and preprocessing
Datasets for training language models fall into two primary categories: general and specialized. General datasets comprise a diverse range of sources, such as web pages, literature, and conversational corpora. For instance, web pages like Wikipedia[52] can contribute to a language model’s broad linguistic understanding. Meanwhile, literary works also serve as a rich reservoir of formal and lengthy texts[53]. These materials are crucial for teaching LLM models complex linguistic constructs, facilitating the modelling of long-range dependencies. Specialized data involves scientific texts and programming-related data. For example, scientific literature comprises a wealth of formal writing imbued with domain-specific knowledge, encompassing academic papers and textbooks. On the other hand, programming data drawn from online question-answering platforms like Stack Exchange [54], along with public software repositories such as GitHub, provide raw material rich with code snippets, comments, and documentation. Incorporating these specialized texts into the training of LLM models can significantly improve LLM’s performance in reasoning and domain-specific knowledge applications. However, before pre-training, it is critical to preprocess the collected datasets, which often contain noisy, redundant, irrelevant, and potentially harmful data. The preprocessing procedure may include quality filtering, de-duplication[49], privacy redaction[55], and tokenization [56].
@@ III-B2 Model training
In the model training process, two pivotal hyperparameters are the batch size and the learning rate. For the pre-training of LLM models, a substantial batch size is required, and recent studies suggest incrementally enlarging the batch size to bolster the stability of the training process [49]. In terms of learning rate adjustments, a widely used strategy is to start with a warm-up phase and then succeed with a cosine decay pattern. This approach helps in achieving a more controllable learning rate schedule. To enhance the training scalability, several key techniques are proposed. For instance, 3D parallelism encompasses data parallelism, pipeline parallelism, and tensor parallelism. Data parallelism involves the replication of the model’s parameters and optimizer states across multiple GPUs[57], allocating to each GPU a subset of data to process and subsequently aggregate the computed gradients. Pipeline parallelism, as detailed in [58], assigns distinct layers of an LLM to various GPUs, allowing the accommodation of larger models within the confines of GPU memory. Tensor parallelism operates on a similar premise by decomposing the tensors[59], especially for the parameter matrices of LLM, facilitating the distribution and computation across multiple GPUs. Meanwhile, ZeRO is also a useful technique [60], which conserves memory by retaining only a portion of the model’s data on each GPU. The remainder of the data is accessible across the GPU network, as needed, effectively addressing memory redundancy concerns.
@@ III-C LLM Fine-tuning
Fine-tuning refers to the process of updating the parameters of pre-trained LLM models to adapt to domain-specific tasks. Although the pre-trained LLM already has vast language knowledge, they lack specialization in specific areas. Fine-tuning overcomes this limitation by allowing the model to learn from domain-specific datasets, making the LLM more effective on specific applications. This subsection will introduce two fine-tuning strategies: instruction and alignment tuning.
@@ III-C1 Instruction tuning
Instruction tuning is a method for fine-tuning pre-trained LLM models using a collection of natural language-formatted instances. This technique aligns closely with supervised fine-tuning and multi-task prompted training, enhancing the LLM’s ability to generalize to unseen tasks, even in multilingual contexts [61]. The process involves collecting or constructing instruction-formatted instances and employing these to fine-tune LLM models in a supervised manner, typically using sequence-to-sequence loss for training. Models like InstructGPT and GPT-4 have demonstrated the effectiveness of instruction tuning in meeting real user needs and improving task generalization [62, 63]. Instruction-formatted instances usually consist of a task description, an optional input, a corresponding output, and possibly a few examples as demonstrations. These instances can originate from various sources, such as traditional NLP task datasets, daily chat data, and synthetic data. Existing research has reformatted traditional NLP datasets with natural language task descriptions to aid LLM models in understanding tasks, proving particularly effective in enhancing task generalization capabilities [61]. The design and quality of instruction instances significantly will impact the model’s performance. Scaling the instructions, for instance, tends to improve generalization ability up to a certain point, beyond which additional tasks may not yield further gains [64]. Diversity in task descriptions and the number of instances per task are also critical, with a smaller number of high-quality instances often sufficing for significant performance improvements [61].
@@ III-C2 Alignment tuning
Alignment tuning aims to ensure LLM models adhere to human values, preventing outputs that could be harmful, biased, or misleading. This concept emerges from the realization that while the LLM excels in various NLP tasks, they may inadvertently generate content that deviates from ethical norms or human expectations [12]. Collecting human feedback is central to the alignment-tuning process. In particular, it involves curating responses from diverse human labellers to guide the LLM toward generating outputs that align with the predefined criteria. Approaches to collecting this feedback include ranking-based methods, where labellers evaluate the quality of model-generated outputs, and question-based methods, where labellers provide insights on specific aspects of the outputs, such as their ethical implications [65].
A prominent technique in alignment tuning is reinforcement learning from human feedback (RLHF), where the model is fine-tuned using reinforcement learning algorithms based on human feedback. This process typically starts with supervised fine-tuning using human-annotated data, followed by training a reward model that reflects human preferences, and finally, fine-tuning the LLM using this reward model. Despite its effectiveness, RLHF can be computationally intensive and complex, necessitating alternative approaches for practical applications [66]. An alternative method for RLHF is direct optimization through supervised learning, which bypasses the complexities of reinforcement learning. This method relies on constructing a high-quality alignment dataset and directly fine-tuning LLM models to adhere to alignment criteria. Although less resource-intensive than RLHF, this approach requires careful dataset construction and may not capture the full range of human values and preferences as effectively as RLHF [67].
@@ III-D LLM Inference and Utilization by Prompting
Prompt engineering is the process in which users design various inputs for AI models to generate desired outputs. Compared with fine-tuning, prompting has no requirements for extra training, producing output instantly based on user inputs. It indicates a straightforward approach to using LLM models, and the rapid response and training-free features make it a promising method for telecom applications. This subsection will introduce key techniques in prompt engineering, including ICL, CoT prompting, and using LLM for complex planning.
@@ III-D1 In-context learning (ICL)
ICL, first introduced with GPT-3 [12], utilizes formatted natural language prompts and integrates task descriptions and examples to guide LLM models in task execution. This approach allows the LLM to recognize and perform new tasks by leveraging contextual information. The design of demonstrations is critical for ICL, encompassing the selection, format, and order of examples. The format of demonstrations involves converting selected examples into a structured prompt, integrating task-specific information and possibly incorporating reasoning enhancements like CoT [64, 68]. The ordering addresses LLM biases, arranging demonstrations based on similarity to the query or employing information-theoretic methods to optimize information conveyance [69, 70].
ICL’s underlying mechanisms include task recognition and task learning, and then LLM models can use pre-trained knowledge and structured prompts to infer and solve new tasks. Task recognition involves LLM models identifying the task type from the provided examples, leveraging pre-existing knowledge from pre-training data [71]. Task learning, on the other hand, refers to LLMs acquiring new task-solving strategies through the given demonstrations, a capability that becomes more pronounced with increasing model size [72]. Recent studies suggest that larger LLM models exhibit an enhanced ability to surpass prior knowledge and learn from the demonstrations provided in ICL settings [73].
@@ III-D2 Chain-of-thought (CoT) prompting
CoT prompting is an advanced strategy to enhance LLM’s performance on complex reasoning tasks, such as arithmetic, commonsense, and symbolic reasoning, by incorporating intermediate reasoning steps into prompts [68]. Differing from ICL’s input-output pairing, CoT prompting enriches prompts with sequences of reasoning steps, guiding LLM models to bridge between questions and answers more effectively. Initially proposed as an ICL extension, CoT augments demonstrations from mere input-output pairs to sequences comprising inputs, intermediate reasoning steps, and outputs [68]. These steps help LLM models navigate complex problem-solving more transparently and logically, though they typically require manual annotation. However, creative phrasings such as “Let’s think step by step” can trigger LLM models to generate CoTs autonomously, which significantly simplifies the CoT implementation.
Despite improvements, CoT prompting faces challenges such as incorrect reasoning and instability. The enhancement strategies include better prompt design by utilizing diverse and complex reasoning paths, advanced generation strategies, and verification-based methods. These methods address generation issues by exploring multiple paths or validating reasoning steps, thus improving result accuracy and stability. Furthermore, extending beyond linear reasoning chains, recent studies propose tree- and graph-structured reasoning to accommodate more complex problem-solving processes[74]. In addition, CoT prompting significantly benefits large-scale LLM models (over 10B parameters) and tasks requiring detailed step-by-step solutions. However, it may underperform in simpler tasks or when traditional prompting is already effective [68].
@@ III-D3 Planning for complex task solving
While ICL and CoT prompting provide a straightforward approach for task solving, they often fall short in complex scenarios like mathematical reasoning and multi-hop question answering [75]. To this end, prompt-based planning has emerged, breaking down intricate tasks into smaller and manageable sub-tasks and outlining action sequences for their resolution. The planning framework for LLM models encompasses three main components: the task planner, the plan executor, and the environment. The task planner devises a comprehensive plan to address the target task, which could be represented as a sequence of actions or an executable program [76]. This plan is then carried out by the plan executor, which can range from text-based models to code interpreters, within an environment that provides feedback on the execution results [77, 74]. In plan generation, LLM models can utilize text-based approaches to produce natural language sequences or code-based methods for generating executable programs, enhancing the verifiability and precision of the planned actions [77]. Feedback acquisition follows, where the LLM evaluates the plan’s efficacy through internal assessments or external signals, refining the strategy based on outcomes from different environments [74]. In addition, the refinement process is crucial for optimizing the plan based on received feedback, and the corresponding methods include reasoning, backtracking, memorization, etc [74].
@@ III-E Evaluation metrics of LLM
Evaluating the performance of LLM models is a multifaceted task and receives increasing attention. This subsection focuses on the evaluation metrics that encompass various dimensions, including accuracy, hallucination, efficiency, and human alignment etc. Each of these aspects plays a crucial role in determining the overall applicability of LLM models in real-world scenarios such as telecom networks. Firstly, accuracy is paramount in evaluating LLM technologies as it directly impacts the model’s reliability and trustworthiness. It measures how well an LLM can understand and process natural language queries, generate relevant and correct responses, and perform specific tasks like translation, summarization, or question-answering. Benchmarks and standardized datasets are often used to quantitatively evaluate the model’s accuracy.
@@
Cloud deployment Cloud deployment is the most straightforward method for LLM deployment. LLM models are usually computationally demanding, and cloud servers can provide abundant computational and storage resources for model training, fine-tuning, and inference. Cloud deployment indicates higher end-to-end latency for implementing user requests, since the inquiries have to be first uploaded and then processed and downloaded. It may prevent the application of some latency-critical applications such as robot control, vehicle-to-vehicle communications, unmanned aerial vehicle (UAV) control, etc.
Network edge deployment Network edge deployment can be an appealing approach to shorten the response time and save backhaul bandwidth to the central cloud. It enables rapid user request processing at edge servers or cloud, achieving shorter end-to-end delay than the central cloud-based approach. Network edge servers are usually resources-constrained, indicating limited computational and storage resources for LLM model fine-tuning and inference. Therefore, some techniques may be exploited, e.g., efficient parameter-efficient fine-tuning[88], split edge learning[89], and quantized training[90]. In addition, model compression is also a promising direction for edge LLM deployment.
On-device deployment On-device LLM is considered a very promising direction to deploy LLM models directly at user devices. It enabled customized LLM models based on specific user requests. Meanwhile, on-device LLMs have the lowest service latency by processing tasks locally. Therefore, it has great potential for implementing real-time tasks. Despite the great advantages, on-device LLMs are still in the very early stages, and the main challenge is to overcome the very limited computational and storage resources at user devices. Apple has proposed a technique to store LLM model parameters on flash memory[91] and achieve a 20 times faster inference speed. Qualcomm also announced a new mobile platform to support popular small-scale LLMs[92]. Therefore, how to utilize limited computation resources to achieve faster inference is the key to on-device LLM deployment.
@@
Cooperative deployment Cooperative deployment is proposed in [93], which involves the interactions between local small models and cloud-based large models. In particular, it assumes that the local model can collect and submit sensor data selectively to the large model, and the large model will update the small-scale local models based on its domain-specific knowledge. The cooperative deployment is a feasible solution to connect small-scale local LLM models to large cloud models. However, the local model updating frequency should also be carefully determined to reduce the burden on cloud LLM models. In addition, note that the inference is still updated locally, and therefore the required computational resources are still challenges. To this end, it may be combined with on-device LLMs to address the resource issues.

### v2

@@ III-B LLM Pre-training
The aim of pre-training language models is to predict the next word within a sentence. After being trained on extensive datasets, LLMs exhibit emergent capabilities in comprehension and reasoning. This subsection will introduce dataset collection, preprocessing, and model training techniques.
@@ III-B1 Dataset collection and preprocessing
Datasets for training language models fall into two primary categories: general and specialized. General datasets comprise a diverse range of sources, such as web pages, literature, and conversational corpora. For instance, web pages like Wikipedia[62] can contribute to a language model’s broad linguistic understanding. Meanwhile, literary works also serve as a rich reservoir of formal and lengthy texts[63]. These materials are crucial for teaching LLMs complex linguistic constructs, facilitating the modelling of long-range dependencies. Specialized data involves scientific texts and programming-related data. For example, scientific literature comprises a wealth of formal writing imbued with domain-specific knowledge, encompassing academic papers and textbooks. On the other hand, programming data drawn from online question-answering platforms like Stack Exchange [64], along with public software repositories such as GitHub, provide raw material rich with code snippets, comments, and documentation. Incorporating these specialized texts into the training of LLMs can significantly improve LLM’s performance in reasoning and domain-specific knowledge applications. However, before pre-training, it is critical to preprocess the collected datasets, which often contain noisy, redundant, irrelevant, and potentially harmful data. The preprocessing procedure may include quality filtering, de-duplication[59], privacy redaction[65], and tokenization [66].
@@ III-B2 Model training
In the model training process, two pivotal hyperparameters are the batch size and the learning rate. For the pre-training of LLMs, a substantial batch size is required, and recent studies suggest incrementally enlarging the batch size to bolster the stability of the training process [59]. In terms of learning rate adjustments, a widely used strategy is to start with a warm-up phase and then succeed with a cosine decay pattern. This approach helps in achieving a more controllable learning rate schedule. To enhance the training scalability, several key techniques are proposed. For instance, 3D parallelism encompasses data parallelism, pipeline parallelism, and tensor parallelism. Data parallelism involves the replication of the model’s parameters and optimizer states across multiple GPUs[67], allocating to each GPU a subset of data to process and subsequently aggregate the computed gradients. Pipeline parallelism, as detailed in [68], assigns distinct layers of an LLM to various GPUs, allowing the accommodation of larger models within the confines of GPU memory. Tensor parallelism operates on a similar premise by decomposing the tensors[69], especially for the parameter matrices of LLM, facilitating the distribution and computation across multiple GPUs. Meanwhile, ZeRO is also a useful technique [70], which conserves memory by retaining only a portion of the model’s data on each GPU. The remainder of the data is accessible across the GPU network, as needed, effectively addressing memory redundancy concerns.
@@ III-C LLM Fine-tuning
Fine-tuning refers to the process of updating the parameters of pre-trained LLMs to adapt to domain-specific tasks. Although the pre-trained LLM already has vast language knowledge, they lack specialization in specific areas. Fine-tuning overcomes this limitation by allowing the model to learn from domain-specific datasets, making the LLM more effective on specific applications. This subsection will introduce two fine-tuning strategies: instruction and alignment tuning.
@@ III-C1 Instruction tuning
Instruction tuning is a method for fine-tuning pre-trained LLMs using a collection of natural language-formatted instances. This technique aligns closely with supervised fine-tuning and multi-task prompted training, enhancing the LLM’s ability to generalize to unseen tasks, even in multilingual contexts [71]. The process involves collecting or constructing instruction-formatted instances and employing these to fine-tune LLMs in a supervised manner, typically using sequence-to-sequence loss for training. Models like InstructGPT and GPT-4 have demonstrated the effectiveness of instruction tuning in meeting real user needs and improving task generalization [72, 73]. Instruction-formatted instances usually consist of a task description, an optional input, a corresponding output, and possibly a few examples as demonstrations. These instances can originate from various sources, such as traditional NLP task datasets, daily chat data, and synthetic data. Existing research has reformatted traditional NLP datasets with natural language task descriptions to aid LLMs in understanding tasks, proving particularly effective in enhancing task generalization capabilities [71]. The design and quality of instruction instances significantly will impact the model’s performance. Scaling the instructions, for instance, tends to improve generalization ability up to a certain point, beyond which additional tasks may not yield further gains [74]. Diversity in task descriptions and the number of instances per task are also critical, with a smaller number of high-quality instances often sufficing for significant performance improvements [71].
@@ III-C2 Alignment tuning
Alignment tuning aims to ensure LLMs adhere to human values, preventing outputs that could be harmful, biased, or misleading. This concept emerges from the realization that while the LLM excels in various NLP tasks, they may inadvertently generate content that deviates from ethical norms or human expectations [58]. Collecting human feedback is central to the alignment-tuning process. In particular, it involves curating responses from diverse human labellers to guide the LLM toward generating outputs that align with the predefined criteria. Approaches to collecting this feedback include ranking-based methods, where labellers evaluate the quality of model-generated outputs, and question-based methods, where labellers provide insights on specific aspects of the outputs, such as their ethical implications [75].
A prominent technique in alignment tuning is reinforcement learning from human feedback (RLHF), where the model is fine-tuned using reinforcement learning algorithms based on human feedback. This process typically starts with supervised fine-tuning using human-annotated data, followed by training a reward model that reflects human preferences, and finally, fine-tuning the LLM using this reward model. Despite its effectiveness, RLHF can be computationally intensive and complex, necessitating alternative approaches for practical applications [76]. An alternative method for RLHF is direct optimization through supervised learning, which bypasses the complexities of reinforcement learning. This method relies on constructing a high-quality alignment dataset and directly fine-tuning LLMs to adhere to alignment criteria. Although less resource-intensive than RLHF, this approach requires careful dataset construction and may not capture the full range of human values and preferences as effectively as RLHF [77]. Additionally, researchers have introduced the direct preference optimization (DPO) technique [78], which eliminates the need for a reward model and allows the model to align directly with preference data. Some advanced LLMs, like Llama 3, utilize both RLHF with proximal policy optimization (PPO) and DPO. However, both RLHF and DPO depend on high-quality human preference data, which is limited and costly to acquire. To address this challenge, methods such as Constitutional AI [79] and RL from AI feedback (RLAIF)  [80] have been developed to generate preference data using LLMs, enabling models to learn from AI feedbacks and facilitating knowledge transfer between models.
@@ III-D LLM Inference and Utilization by Prompting
Prompt engineering is the process in which users design various inputs for AI models to generate desired outputs. Compared with fine-tuning, prompting has no requirements for extra training, producing output instantly based on user inputs. It indicates a straightforward approach to using LLMs, and the rapid response and training-free features make it a promising method for telecom applications. This subsection will introduce key techniques in prompt engineering, including ICL, CoT prompting, LLM for complex planning, and self-refinement with iterative feedback. The comparisons among different prompt engineering techniques are shown in Fig. 3.
Figure 3: Comparison of various prompt engineering techniques [11, 81]. Different from task-specific examples in ICL demonstrations, CoT prompting additionally incorporates intermediate reasoning steps in demonstrations. Prompt-based planning breaks down intricate tasks into manageable sub-tasks and outlines action sequences for their resolution. Self-refine enhances LLM’s outputs through iterative feedback and refinement.
@@ III-D1 In-context learning (ICL)
ICL, first introduced with GPT-3 [58], utilizes formatted natural language prompts and integrates task descriptions and examples to guide LLMs in task execution. This approach allows the LLM to recognize and perform new tasks by leveraging contextual information. The design of demonstrations is critical for ICL, encompassing the selection, format, and order of examples. The format of demonstrations involves converting selected examples into a structured prompt, integrating task-specific information and possibly incorporating reasoning enhancements like CoT [74, 82]. The ordering addresses LLM biases, arranging demonstrations based on similarity to the query or employing information-theoretic methods to optimize information conveyance [83, 84]. ICL’s underlying mechanisms include task recognition and task learning, and then LLMs can use pre-trained knowledge and structured prompts to infer and solve new tasks. Task recognition involves LLMs identifying the task type from the provided examples, leveraging pre-existing knowledge from pre-training data [85]. Task learning, on the other hand, refers to LLMs acquiring new task-solving strategies through the given demonstrations, a capability that becomes more pronounced with increasing model size [86]. Recent studies suggest that larger LLMs exhibit an enhanced ability to surpass prior knowledge and learn from the demonstrations provided in ICL settings [87].
@@ III-D2 Chain-of-thought (CoT) prompting
CoT prompting is an advanced strategy to enhance LLM’s performance on complex reasoning tasks, such as arithmetic, commonsense, and symbolic reasoning, by incorporating intermediate reasoning steps into prompts [82]. Differing from ICL’s input-output pairing, CoT prompting enriches prompts with sequences of reasoning steps, guiding LLMs to bridge between questions and answers more effectively. Initially proposed as an ICL extension, CoT augments demonstrations from mere input-output pairs to sequences comprising inputs, intermediate reasoning steps, and outputs [82]. These steps help LLMs navigate complex problem-solving more transparently and logically, though they typically require manual annotation. However, creative phrasings such as “Let’s think step by step” can trigger LLMs to generate CoTs autonomously, which significantly simplifies the CoT implementation.
Despite improvements, CoT prompting faces challenges such as incorrect reasoning and instability. The enhancement strategies include better prompt design by utilizing diverse and complex reasoning paths, advanced generation strategies, and verification-based methods. These methods address generation issues by exploring multiple paths or validating reasoning steps, thus improving result accuracy and stability. Furthermore, extending beyond linear reasoning chains, recent studies propose tree- and graph-structured reasoning to accommodate more complex problem-solving processes[88]. In addition, CoT prompting significantly benefits large-scale LLMs (over 10B parameters) and tasks requiring detailed step-by-step solutions. However, it may underperform in simpler tasks or when traditional prompting is already effective [82].
@@ III-D3 Planning for complex task solving
While ICL and CoT prompting provide a straightforward approach for task solving, they often fall short in complex scenarios like mathematical reasoning and multi-hop question answering [89]. To this end, prompt-based planning has emerged, breaking down intricate tasks into smaller and manageable sub-tasks and outlining action sequences for their resolution. The planning framework for LLMs encompasses three main components: the task planner, the plan executor, and the environment. The task planner devises a comprehensive plan to address the target task, which could be represented as a sequence of actions or an executable program [90]. This plan is then carried out by the plan executor, which can range from text-based models to code interpreters, within an environment that provides feedback on the execution results [91, 88]. In plan generation, LLMs can utilize text-based approaches to produce natural language sequences or code-based methods for generating executable programs, enhancing the verifiability and precision of the planned actions [91]. Feedback acquisition follows, where the LLM evaluates the plan’s efficacy through internal assessments or external signals, refining the strategy based on outcomes from different environments [88]. In addition, the refinement process is crucial for optimizing the plan based on received feedback, and the corresponding methods include reasoning, backtracking, memorization, etc [88].
III-D4 Self-refinement with iterative feedback
Considering that LLMs may not generate correct answers initially, self-refine has recently emerged to improve their outputs through iterative feedback and refinement. Among these studies, Madaan et al. [81] first use an LLM to generate initial outputs, and then employ the same LLM to provide specific feedback on these outputs. Note that this feedback is actionable, containing concrete steps to further improve the initial outputs. With such specific and actionable feedback, the same LLM can iteratively refine its outputs until performance converges. Furthermore, Hu et al. [92] propose a self-refined LLM (named TrafficLLM) specifically designed for communication traffic prediction, which leverages in-context learning to enhance predictions through a three-step process: traffic prediction, feedback generation, and prediction refinement. Following the comprehensive feedback, refinement demonstration prompts enable the same LLM to refine its performance on target tasks.
@@ III-E Evaluation metrics of LLM
Evaluating the performance of LLMs is a multifaceted task and receives increasing attention. This subsection focuses on the evaluation metrics that encompass various dimensions, including accuracy, hallucination, efficiency, and human alignment etc. Each of these aspects plays a crucial role in determining the overall applicability of LLMs in real-world scenarios such as telecom networks. Firstly, accuracy is paramount in evaluating LLM technologies as it directly impacts the model’s reliability and trustworthiness. It measures how well an LLM can understand and process natural language queries, generate relevant and correct responses, and perform specific tasks like translation, summarization, or question-answering. Benchmarks and standardized datasets are often used to quantitatively evaluate the model’s accuracy.
@@
Cloud deployment Cloud deployment is the most straightforward method for LLM deployment. LLMs are usually computationally demanding, and cloud servers can provide abundant computational and storage resources for model training, fine-tuning, and inference. Cloud deployment indicates higher end-to-end latency for implementing user requests, since the inquiries have to be first uploaded and then processed and downloaded. It may prevent the application of some latency-critical applications such as robot control, vehicle-to-vehicle communications, unmanned aerial vehicle (UAV) control, etc.
Network edge deployment Network edge deployment can be an appealing approach to shorten the response time and save backhaul bandwidth to the central cloud. It enables rapid user request processing at edge servers or cloud, achieving shorter end-to-end delay than the central cloud-based approach. Network edge servers are usually resources-constrained, indicating limited computational and storage resources for LLM fine-tuning and inference. Therefore, some techniques may be exploited, e.g., efficient parameter-efficient fine-tuning[103], split edge learning[104], and quantized training[105]. In addition, model compression is also a promising direction for edge LLM deployment.
On-device deployment On-device LLM is considered a very promising direction to deploy LLMs directly at user devices. It enabled customized LLMs based on specific user requests. Meanwhile, on-device LLMs have the lowest service latency by processing tasks locally. Therefore, it has great potential for implementing real-time tasks. Despite the great advantages, on-device LLMs are still in the very early stages, and the main challenge is to overcome the very limited computational and storage resources at user devices. Apple has proposed a technique to store LLM parameters on flash memory[106] and achieve a 20 times faster inference speed. Qualcomm also announced a new mobile platform to support popular small-scale LLMs[107]. Therefore, how to utilize limited computation resources to achieve faster inference is the key to on-device LLM deployment.
@@
Cooperative deployment Cooperative deployment is proposed in [108], which involves the interactions between local small models and cloud-based large models. In particular, it assumes that the local model can collect and submit sensor data selectively to the large model, and the large model will update the small-scale local models based on its domain-specific knowledge. The cooperative deployment is a feasible solution to connect small-scale local LLMs to large cloud models. However, the local model updating frequency should also be carefully determined to reduce the burden on cloud LLMs. In addition, note that the inference is still updated locally, and therefore the required computational resources are still challenges. To this end, it may be combined with on-device LLMs to address the resource issues.